OpenStudy (anonymous):

SAT: The difference of two numbers is 4 and their sum is -7. What is their product?

OpenStudy (anonymous):

Let the numbers be x and y therefore x - y = 4 and x + y = -7 Adding the two equations we get 2x = -3 i.e. x =-3/2 = -1.5 Therefore y = -5.5 their product i.e. xy = (-1.5) (-.5.5) = 8.25 Thus their product is 8.25

ganeshie8 (ganeshie8):

another way :- \( \large (x-y)^2 = (x+y)^2 - 4xy \) \( \large 4^2 = (-7)^2 - 4xy \)

ganeshie8 (ganeshie8):

you can solve xy